You are considering purchasing shares of a publicly traded company. The most recent dividend paid by the company was $6.62. The appropriate equity cost of capital for the company is 7.9%. The company is experiencing growth, and the expected growth rates for the next 4 years are listed below. After this period of increasing growth passes, the company expects growth to stabilize at a long run growth rate of 3.0%.

Year 1 = 1.5% Growth Rate

Year 2 = 3.0% Growth Rate

Year 3 = 4.5% Growth Rate

Year 4 = 6% Growth Rate

What are the dividends necessary? (5 dividends) How to find these dividends is needed.
Answer #1

Most recent dividend D0 =6.62
Dividend year 1 (D1) =D0*(1+Year 1 growth rate) =6.62*(1+1.5%) =6.7193
Dividend year 2 (D2) =D1*(1+Year 2 growth rate) =6.7193*(1+3%) =6.920879 or 6.9209
Dividend year 3 (D3) =D2*(1+Year 3 growth rate) =6.920879*(1+4.5%) =7.232318555 or 7.2323
Dividend year 4 (D4) =D3*(1+Year 4 growth rate) =7.232318555*(1+6%) =7.6662576683 or 7.6663
Dividend year 5 (D5) =D4*(1+Year 5 growth rate) =7.6662576683*(1+3%) =7.896245398349 or 7.8962